The price elasticity of demand for chicken is estimated to be 0.45 if the price of chicken increased by 08 percent what will be the expected percentage decreased in the quality of chicken sold
1
Expert's answer
2020-06-25T11:04:38-0400

Price elasticity of demand (PED) = 0.45

Increase in Pricechicken by 8%

PED = "\frac {\% change in Quantity} {\% change in Price}"

PED = "\frac{∆Q/Q} {∆P/P}"

% change in Quantity = "PED \times \% change in Price"

% change in Quantity = "0.45\times8\% = 3.6\%"

The quantity demanded of chicken will decrease by 3.6%.The price elasticity of demand for chicken (0.45) is inelastic since it is less than 1 and as such, the huge increase in the price of chicken by 8 percent has a small effect in the change in quantity of chicken. It only causes a decrease in the quantity of chicken consumed by 3.6%.